# TDDSR: Single-Step Diffusion with Two Discriminators for Super Resolution

### v1

@@ Abstract
Super-resolution methods are increasingly being specialized for both real-world and face-specific tasks. However, many existing approaches rely on simplistic degradation models, which limits their ability to handle complex and unknown degradation patterns effectively. While diffusion-based super-resolution techniques have recently shown impressive results, they are still constrained by the need for numerous inference steps. To address this, we propose TDDSR, an efficient single-step diffusion-based super-resolution method. Our method, distilled from a pre-trained teacher model and based on a diffusion network, performs super-resolution in a single step. It integrates a learnable downsampler to capture diverse degradation patterns and employs two discriminators—one for high-resolution and one for low-resolution images—to enhance the overall performance. Experimental results demonstrate its effectiveness across real-world and face-specific SR tasks, achieving performance comparable to, or even surpassing, another single-step method, previous state-of-the-art models, and the teacher model.
(a) LR input
(b) ResShift-15
(c) SinSR-1
(d) Ours-1
@@ 1 Introduction
The quality of images in real world degrades from various factors, including blurs, compression artifacts, color inaccuracies, and sensor noises. A major challenge in Super-Resolution (SR) is that these diverse degradation patterns are often unknown. Traditionally, degradations have been assumed and modeled using a simple process like Gaussian noise or bicubic downsampling. More sophisticated degradation models [61, 52, 62, 36] have been proposed to reflect real-world conditions better. While these models can be managed effectively, extending their application to unforeseen degradations remains challenging, with complications in achieving realistic image reconstruction.
@@
Building upon these developments, we propose a novel approach to Super-Resolution that leverages the strengths of both diffusion models and GANs while considering diverse degradation patterns. Specifically, we introduce a single-step diffusion model for image super-resolution, distilled from a pre-trained diffusion model and augmented by two discriminators. Our methodology involves employing a deterministic sampling strategy and distilling the deterministic generation function [54] into a student network. To further enhance performance, we incorporate two discriminators: (i) an HR discriminator, which utilizes ground-truth high-resolution (HR) images to enhance overall image quality, and (ii) a Downsampler & LR discriminator, which uses a flexible downsampler to provide detailed degradation information and compares the generated output with low-resolution (LR) images.
Our contributions are summarized as follows:
  • We propose a novel approach combining the diffusion network with two discriminators.
  • The method achieves on-par state-of-the-art results with a single-step process.
  • Performance is enhanced without increasing the complexity of the student network, keeping the number of parameters constant.
  • We demonstrate the excellent performance of the method in real-world SR tasks and on face data.
@@ 2.2 Image Super-Resolution
Generative Adversarial Networks (GANs)[10] consists of two competing networks: a generator and a discriminator. The goal of the generator is to synthesize high-resolution images that appear highly realistic, while the discriminator strives to distinguish between real images and those generated by the generator. This adversarial training framework has been effectively employed in the Image Super-Resolution (ISR) task [23, 53, 52, 61, 27, 16]. In particular, GANs have been utilized to achieve realistic ISR, beginning with SRGAN[23], which assumed bicubic downsampling for image degradation. Subsequent research has explored more complex degradation models, such as randomly shuffled and high-order processes, as illustrated by BSRGAN[61] and Real-ESRGAN[52]. These advancements have shown the potential of GAN-based models to handle the complexities of real-world image super-resolution, moving beyond simple degradation models and enhancing the perceptual quality of the generated images.
@@
Using Downsampler.
When using a learnable downsampler for super-resolution, it typically involves generating low-resolution (LR) images. One approach [3] employs a simple downsampling network to create paired LR-HR images for training with GAN. Another method [1, 51] focuses on learning degradation features or representations to enhance specific tasks. [35] introduces downscaling loss, explicitly penalizing a proposed super-resolved image for deviating from its LR input. Meanwhile, [9] proposes a learnable downsampler that directly influences the upsampler, closely aligning with our approach.
@@ 2.3 Accelerating Diffusion Sampling
Figure 2: The overall framework of TDDSR. The student network $f_{\phi}$ is trained to learn a deterministic mapping from $x_{T}$ to $\hat{x}_{0}$ in just one step, guided by a pre-trained teacher network $f_{\theta}$. The student’s output $f_{\phi}(x_{T},y,T)$ then goes to the High-Resolution Discriminator $\mathcal{D}_{H}$. Simultaneously, $f_{\phi}(x_{T},y,T)$ is also processed by a Learnable Downsampler $G$ and Low-Resolution Discriminator $\mathcal{D}_{L}$.
@@ 4 Methodology
Our training process, as shown in Fig. 2, involves three paths. The first path focuses on distilling the knowledge from the teacher model into the student model, enabling the Super-Resolution (SR) task to be performed in a single step. The second path involves passing the output $f_{\phi}(x_{T},y,T)$ generated by the student model, along with the ground-truth high-resolution (HR) image, to the HR discriminator, which distinguishes between the two. In the third path, the student’s output $f_{\phi}(x_{T},y,T)$ is processed through a downsampler to estimate a low-resolution image $\hat{y}$. This estimated image $\hat{y}$ is then compared to the original low-resolution image $y$ by the LR discriminator, which determines whether the image is real or fake. We drew inspiration from DiscoGAN[22] and CycleGAN[66], particularly in their use of cycle consistency loss for bidirectional image translation. These models learn mappings between two domains, $x$ and $y$, where a function maps $x$ to $y$, and a reverse function maps $y$ back to $x$. The cycle consistency loss ensures accurate image reconstruction after both transformations. Similarly, we apply this concept to super-resolution by augmenting the student network, which transforms LR to HR, with a learnable downsampler that performs the HR to LR mapping. Two discriminators compare the generated images with their ground-truths, enforcing consistency and fidelity across both resolutions.
@@ 4.2 High-Resolution Image discriminator
Methods Datasets
RealSR RealSet65
CLIPIQA(↑) MUSIQ(↑) CLIPIQA(↑) MUSIQ(↑)
RankSRGAN 0.5821 62.0984 0.5600 51.8125
[…]
ResShift-15 0.6014 58.6475 0.6489 60.7733
[…]
Metrics
Methods PSNR(↑) SSIM(↑) LPIPS(↓) CLIPIQA(↑) MUSIQ(↑)
RankSRGAN 26.51 0.7526 0.1217 0.6402 64.6863
[…]
ResShift-15 27.2753 0.7626 0.2005 0.6729 65.5703
[…]
@@ 4.3 Learnable Downsampler and Low-Resolution Image Discriminator
The key distinction between prior works and our method lies in its backbone architecture. Leveraging a diffusion network, our approach integrates a flexible downsampler that spatially adapts the downsampling kernel, enabling it to effectively capture diverse degradation patterns. This design enhances the performance of the diffusion network by providing richer and more varied information during training. The Learnable Downsampler is composed of convolutional layers designed to dynamically adjust different degradation patterns, offering the adaptability required to address a wide range of image degradation scenarios.
@@
where $y$ is the ground-truth low-resolution image, and $\hat{y}=G(f_{\phi}(x_{T},y,T))$ is the LR image predicted by the learnable downsampler $G$. The downsampler’s adversarial objective amounts to
@@ 5.1 Real-world Super-Resolution
Results.
The results on RealSR and RealSet65 are presented in Table 1. The table demonstrates that the proposed approach outperforms SinSR and the teacher model, even with only a single inference step. In terms of the CLIPIQA metric, TDDSR achieves the best performance across both datasets. The results of the DIV2K validation dataset are shown in Table 2. While the PSNR and SSIM values decrease compared to the teacher model due to the reduction in inference steps from 15 to 1, our one-step model demonstrates superior performance in perceptual quality metrics, including LPIPS and the state-of-the-art metrics CLIPIQA and MUSIQ. This indicates that our proposed one-step model performs well from a perceptual and realism perspective. Additionally, compared to SinSR, which also employs a single inference step, our model achieves better performance in all metrics except CLIPIQA. Figure 3 presents some visual comparisons among the teacher model ResShift, the single-step model SinSR, and our proposed single-step model. Besides, more comparisons are in the supplementary material.
5.2 Face Super-Resolution
(a) LR input
(b) Ours
(c) GT
@@ 5.2 Face Super-Resolution
Results.
In Table 4, we present the quantitative results of 64$\times$64 images by $\times$4 SR, demonstrating that our method achieves superior performance in terms of PSNR, SSIM, and FID metrics compared to other approaches. Qualitative comparisons of the LR input, our SR results, and the ground-truth HR images are provided in Figure 5. We also present the results of 1024$\times$1024 output images in Table 3, where both our method and SinSR used models trained on 256$\times$256 images. Despite this, our method achieves the best performance in terms of FID and ranks second in both PSNR and MS-SSIM. Visual comparisons, depicted in Figure 4, highlight that our method effectively captures fine details such as wrinkles and skin smoothness, closely approximating the quality of GT images. We directly compared our proposed method with the ground-truth to evaluate how closely our results align with the ideal reference images, thus providing a deeper understanding of the perceptual quality and fidelity of our SR images. Additionally, further comparisons can be found in the supplementary material.
Metrics
Methods PSNR(↑) SSIM(↑) MS-SSIM(↑) FID(↓)
SRGAN 21.49 0.515 0.807 60.67
ESRGAN 19.84 0.353 0.782 72.73
[…]
SRFBN 27.90 0.822 0.931 17.14
CAGFace 34.10 0.906 0.971 12.40
[…]
@@ Results.
5.3 Ablation Study
We conducted an ablation study on combining three loss paths, as shown in Table 5. The study compares the following losses: the distillation loss (Distill) described in Sec 4.1, the HR loss using the student’s SR images with an HR discriminator as explained in Sec 4.2, and the LR loss using a learnable downsampler with an LR discriminator, also described in Sec 4.3. We evaluated the performance on real-world datasets, RealSR and RealSet65, using CLIPIQA and MUSIQ metrics. The results indicate that all three losses are essential. While combining the distillation loss with either the HR loss or the LR loss improved performance, the best improvement was achieved when all three losses were combined.
@@ 5.3 Ablation Study
6 Conclusion
In this work, we introduced a single-step diffusion model equipped with two discriminators to enhance inference efficiency while preserving high generative performance. Our method uniquely updates the student network cyclically, considering both HR and LR perspectives, and integrates an adversarial loss via the discriminators. Additionally, we introduced a learnable downsampler to better capture diverse degradation patterns. We demonstrated the effectiveness of our approach on both Real-ISR and Face-SR tasks. While our method shows improved performance compared to the prior one-step approach, there remain still challenges, particularly with fine-scale details such as very small scene text. We also anticipate that training on larger datasets could further improve the model’s overall generative capabilities.
@@ 6 Conclusion
In this Supplementary Material, we provide additional qualitative results for both real-world and facial datasets, as well as further ablation studies.
S.1. Experiment Details
@@
Similar to the main paper, we present experimental results for both Real-World Super-Resolution and Face Super-Resolution, along with visual comparisons of the outcomes. Fig 6 illustrates the results of 4$\times$ super-resolution on real-world datasets, comparing our method to ResShift (multi-step)  [60] and SinSR (single-step) [54]. Fig 7 and Fig 8 provide detailed comparisons for face datasets with 1024$\times$1024 output resolutions. Our model was trained on ground truth images of size 256, and while the current results are promising relative to the ground truth—demonstrating the scalability of the model—we believe that training the model with 1024$\times$1024 ground truth data could further enhance both quantitative and qualitative performance. The comparisons are highlighted by observing the boxed areas in the low-resolution (LR) input, where the colored boxes correspond to our results and the ground truth. Fig 9 and Fig 10 show comparisons using a 64$\times$64 input and a 256$\times$256 output. Furthermore, when calculating the FID [13] score, we used the Inception-v3 pool3 layer to ensure consistency with the models and results reported in previous studies [15].
S.2. Ablation Study
Three Losses.
In the main paper, we conducted an ablation study on three losses: Distillation loss, HR loss (using an HR discriminator), and LR loss (using a learnable downsampler and LR discriminator). From a visual perspective, as shown in Fig 11, when zoomed in for detailed inspection, there is a noticeable difference, the combination of all three losses helping to reduce artifacts.
Pretrained Discriminators.
We also conducted an ablation study to compare the results with and without pretraining the two discriminators. The HR discriminator was pre-trained using fake images generated by the Teacher model, ResShift [60]. For the LR discriminator, we employed a learnable downsampler to generate fake LR images by training on both HR and LR ground-truth images.
The results, based on experiments using the DIV2K validation dataset, are presented in Table 6. In these experiments, v1 corresponds to the model without pre-trained discriminators, while v2 corresponds to the model with pre-trained discriminators. The use of pre-trained discriminators yields better performance on the CLIPIQA and MUSIQ metrics. However, for PSNR, SSIM, and LPIPS, the model without pre-trained discriminators produces better results. Given that the differences in performance are not substantial, we chose not to pre-train the discriminators in this study.
Methods Metrics
PSNR(↑) SSIM(↑) LPIPS(↓) CLIPIQA(↑) MUSIQ(↑)
Ours v1 26.8800 0.7466 0.1977 0.6855 65.8653
Ours v2 26.8058 0.7413 0.2023 0.6971 66.5178
Table 6: Ablation study comparing the results with and without pretraining the two discriminators

### v2

@@ Abstract
Super-resolution methods are increasingly becoming popular for both real-world and face-specific tasks. Many existing approaches, however, rely on simplistic degradation models, which limits their ability to handle complex and unknown degradation patterns effectively. While diffusion-based super-resolution techniques have recently shown impressive results, they are still constrained by the need for numerous inference steps. To address this, we propose TDDSR, an efficient single-step diffusion-based super-resolution method. Our method, distilled from a pre-trained teacher model and based on a diffusion network, performs super-resolution in a single step. It integrates a learnable diffusion-based downsampler to capture diverse degradation patterns and employs two discriminators—one for high-resolution and one for low-resolution images—to enhance the overall performance. Experimental results demonstrate its effectiveness across real-world and face-specific SR tasks, achieving performance beyond other state-of-the-art models and comparable to previous diffusion methods with multiple sampling steps.
@@ 1 Introduction
The quality of images in the real world degrades from various factors, including blurs, compression artifacts, color inaccuracies, and sensor noises. A major challenge in Super-Resolution (SR) is that these diverse degradation patterns are often unknown. Traditionally, degradations have been assumed and modeled using a simple process like Gaussian noise or bicubic downsampling. More sophisticated degradation models [61, 52, 62, 36] have been proposed to reflect real-world conditions better. While these models can be managed effectively, extending their application to unforeseen degradations remains challenging, with complications in achieving realistic image reconstruction.
@@
Figure 1: Both the student network (low-to-high) and the downsampler (high-to-low) are diffusion-based architectures. In the latent space, the output of the student network is conditioned and fed into the downsampler. The output of the downsampler $\hat{y}$ is then compared with the original low-resolution $y$ during training.
@@
Building upon these developments, we propose a novel approach to Super-Resolution that leverages the strengths of both diffusion models and GANs while considering diverse degradation patterns. Specifically, we introduce a single-step diffusion model for image super-resolution, distilled from a pre-trained diffusion model and augmented by two discriminators. Our methodology involves employing a deterministic sampling strategy and distilling the deterministic generation function [54] into a student network. To further enhance performance, we incorporate two discriminators: (i) an HR discriminator, which utilizes ground-truth high-resolution (HR) images to enhance overall image quality, and (ii) a Downsampler & LR discriminator, which uses a flexible diffusion-based downsampler to provide detailed degradation information and compares the generated output with low-resolution (LR) images as shown in Fig. 1. Our contributions are summarized as follows:
  • We propose a novel approach combining the diffusion networks with two discriminators.
  • Both the upsampler and downsampler are diffusion networks learnt via their interaction in the latent space.
  • The method achieves state-of-the-art results with a single-step process. Performance is enhanced without increasing the complexity of the student network, keeping the number of parameters constant.
  • We demonstrate the excellent performance of the method in real-world SR tasks and face data.
@@ 2.2 Image Super-Resolution
Generative Adversarial Networks (GANs)[10] consist of two competing networks: a generator and a discriminator. The goal of the generator is to synthesize high-resolution images that appear highly realistic, while the discriminator strives to distinguish between real images and those generated by the generator. This adversarial training framework has been effectively employed in the Image Super-Resolution (ISR) task [23, 53, 52, 61, 27, 16]. In particular, GANs have been utilized to achieve realistic ISR, beginning with SRGAN[23], which assumed bicubic downsampling for image degradation. Subsequent research has explored more complex degradation models, such as randomly shuffled and high-order processes, as illustrated by BSRGAN[61] and Real-ESRGAN[52]. These advancements have shown the potential of GAN-based models to handle the complexities of real-world image super-resolution, moving beyond simple degradation models and enhancing the perceptual quality of the generated images.
Figure 2: The overall framework of TDDSR. The student network $f_{\phi}$ is trained to learn a deterministic mapping from $x_{T}$ to $\hat{x}_{0}$ in just one step, guided by a pre-trained teacher network $f_{\theta}$. The student’s output $f_{\phi}(x_{T},y,T)$ then goes to the High-Resolution Discriminator $\mathcal{D}_{H}$. Simultaneously, $f_{\phi}(x_{T},y,T)$ is jointly learnt with a Learnable Downsampler $G$ and Low-Resolution Discriminator $\mathcal{D}_{L}$ in end-to-end fashion.
@@
Using a Downsampler.
When using a learnable downsampler for super-resolution, it typically involves generating low-resolution (LR) images. One approach [3] employs a simple downsampling network to create paired LR-HR images for training with GAN. Another method [1, 51] focuses on learning degradation features or representations to enhance specific tasks. [35] introduces a downscaling loss, explicitly penalizing super-resolved images that deviate from LR inputs. Meanwhile, [9] proposes a learnable downsampler that directly influences the upsampler, closely aligning with our approach.
@@ 2.3 Accelerating Diffusion Sampling
(a) LR input
(b) ResShift-15
(c) SinSR-1
(d) Ours-1
@@ 4 Methodology
Our training process, as shown in Fig. 2, involves three paths. The first path focuses on distilling the knowledge from the teacher model into the student model, enabling the Super-Resolution (SR) task to be performed in a single step. The second path involves passing the output $f_{\phi}(x_{T},y,T)$ generated by the student model, along with the ground-truth high-resolution (HR) image, to the HR discriminator, which distinguishes between the two. In the third path, the student’s output $f_{\phi}(x_{T},y,T)$ is processed through a downsampler to estimate a low-resolution image $\hat{y}$. This estimated image $\hat{y}$ is then compared to the original low-resolution image $y$ by the LR discriminator, which determines whether the image is real or fake.
We drew inspiration from DiscoGAN [22] and CycleGAN [66], particularly in their use of cycle consistency loss for bidirectional image translation. These models learn mappings between two domains, $x$ and $y$, where a function maps $x$ to $y$, and a reverse function maps $y$ back to $x$. The cycle consistency loss ensures accurate image reconstruction after both transformations. Similarly, we apply this concept to super-resolution by augmenting the student network, which transforms LR to HR, with a learnable downsampler that performs the HR to LR mapping. Two discriminators compare the generated images with their ground-truths, enforcing consistency and fidelity across both resolutions.
@@ 4.3 Learnable Downsampler and Low-Resolution Image Discriminator
The key distinction between prior works and our method lies in its backbone architecture. Leveraging a diffusion network, our approach integrates a flexible downsampler that spatially adapts the downsampling kernel, enabling it to effectively capture diverse degradation patterns. This design enhances the performance of the diffusion network by providing richer and more varied information during training. The Learnable Downsampler utilizes a diffusion network with the same structure as the student network, designed to dynamically adjust different degradation patterns, offering the adaptability required to address a wide range of image degradation scenarios. As both are diffusion-based networks, the output of the student network is concatenated with noise in the latent space and then fed as input to the downsampler. See Fig. 1.
@@
where $y$ is the ground-truth low-resolution image, and $\hat{y}$ is the LR image predicted by the learnable downsampler $G$. The downsampler’s adversarial objective amounts to
@@
4.4 Losses for Two Diffusion Networks
Both the student network (low-to-high) and the downsampler network (high-to-low) employ diffusion models, with their respective total losses defined as follows. With $\mathcal{L}_{denoise}=||\epsilon-\epsilon_{\theta}(x_{t})||^{2}_{2}$ and $\theta$ representing the model parameters, the total loss of the student network is defined as follows
$\mathcal{L}_{student}$ $=\mathcal{L}_{denoise}+\mathcal{L}_{distill}+\mathcal{L}^{G}_{H}+\mathcal{L}^{G}_{L}.$ (11)
The total loss of the downsampler is defined as follows
$\mathcal{L}_{downsampler}$ $=\mathcal{L}_{denoise}+\mathcal{L}^{G}_{L}.$ (12)
@@ 5.1 Real-world Super-Resolution
Results.
The results on RealSR and RealSet65 are presented in Table 1. The table demonstrates that the proposed approach outperforms SinSR and the teacher model, even with only a single inference step. In terms of both the CLIPIQA and MUISQ metrics, TDDSR achieves the best performance across both datasets. The results of the DIV2K validation dataset are shown in Table 2. While the PSNR and SSIM values decrease compared to the teacher model due to the reduction in inference steps from 15 to 1, our one-step model demonstrates superior performance in perceptual quality metrics, including LPIPS and the state-of-the-art metrics CLIPIQA and MUSIQ. This indicates that our proposed one-step model performs well from a perceptual and realism perspective. Additionally, compared to SinSR, which also employs a single inference step, our model achieves better performance in all metrics except CLIPIQA. Figure 4 presents some visual comparisons among the teacher model ResShift, the single-step model SinSR, and our proposed single-step model. Besides, more comparisons are in the supplementary material.
Datasets
RealSR RealSet65
Methods CLIPIQA(↑) MUSIQ(↑) CLIPIQA(↑) MUSIQ(↑)
[…]
ResShift (15 sampling steps) 0.6014 58.6475 0.6489 60.7723
[…]
Ours (single step) 0.7240 63.2628 0.7428 64.0630
[…]
Metrics
Methods PSNR(↑) SSIM(↑) LPIPS(↓) CLIPIQA(↑) MUSIQ(↑)
RankSRGAN 26.5100 0.7526 0.1217 0.6402 64.6863
[…]
ResShift (15 sampling steps) 27.2753 0.7626 0.2005 0.6729 65.5703
[…]
5.2 Face Super-Resolution
@@
Results.
In Table 3, we present the quantitative results of 64$\times$64 images by $\times$4 SR, demonstrating that our method achieves superior performance in terms of PSNR, SSIM, and FID metrics compared to other approaches. Qualitative comparisons of the LR input, our SR results, and the ground-truth HR images are provided in Figure 7. We also present results on 1024$\times$1024 output images, using a model trained on 256$\times$256 images, to demonstrate scalability. Visual comparisons in Figure 5 show that our method effectively captures fine details, such as wrinkles and skin smoothness, closely approximating the quality of the GT images. We directly compared our proposed method with the ground-truth to evaluate how closely our results align with the ideal reference images, thus providing a deeper understanding of the perceptual quality and fidelity of our SR images. Additionally, further comparisons can be found in the supplementary material.
(a) LR input
(b) Ours
(c) GT
@@ Results.
5.3 Ablation Study
We conducted an ablation study on combining three loss paths, as shown in Table 4. The study compares the following losses: the distillation loss (Distill) described in Sec 4.1, the HR loss using the student’s SR images with an HR discriminator as explained in Sec 4.2, and the LR loss using a learnable downsampler with an LR discriminator, also described in Sec 4.3. For this study, the downsampler was implemented using convolutional layers. We evaluated the performance on real-world datasets, RealSR and RealSet65, using CLIPIQA and MUSIQ metrics. The results indicate that all three losses are essential. While combining the distillation loss with either the HR loss or the LR loss improved performance, the best improvement was achieved when all three losses were combined.
6 Conclusions
In this work, we introduced a single-step diffusion model equipped with two discriminators to enhance inference efficiency while preserving high generative performance. Our method uniquely updates the student network cyclically, considering both HR and LR perspectives, and integrates an adversarial loss via the discriminators. Additionally, we introduced a learnable diffusion-based downsampler to better capture diverse degradation patterns. We demonstrated the effectiveness of our approach on both Real-ISR and Face-SR tasks. While our method shows improved performance compared to the prior one-step approach, there remain still challenges, particularly with fine-scale details such as very small scene text. We also anticipate that training on larger datasets could further improve the model’s overall generative capabilities.
@@ 6 Conclusions
In this supplementary material, we present additional qualitative results for both real-world and facial datasets, along with further ablation studies. We provide additional mathematical explanations for the learnable downsampler and demonstrate the effects of using a diffusion-based architecture for the downsampler.
1 Experiment Details
1.1 Qualitative Results
@@
Similar to the main paper, we present experimental results for both Real-World Super-Resolution and Face Super-Resolution, along with visual comparisons of the outcomes. Fig 8 illustrates the results of 4$\times$ super-resolution on real-world datasets, comparing our method (single-step) to ResShift (multi-step)  [60] and SinSR (single-step) [54]. Fig 9 and Fig 10 provide detailed comparisons for face datasets with 1024$\times$1024 output resolutions. Our model, trained on ground-truth images of size 256, demonstrates strong scalability. The comparisons are highlighted by observing the boxed areas in the low-resolution (LR) input, where the colored boxes correspond to our results and the ground truth. Fig 12 and Fig 14 show comparisons using a 64$\times$64 input and a 256$\times$256 output. The results for SinSR are shown after training to the same number of epochs as our model. Furthermore, when calculating the FID [13] score, we used the Inception-v3 pool3 layer to ensure consistency with the models and results reported in previous studies [15].
1.2 Learnable Downsampler
The process of interaction between the Student network(low-to-high) and the Downsampler network(high-to-low) within the latent space is formalized as follows. The initial input to the Student network is a low-resolution image $y$ processed by an encoder and then conditioned on Gaussian noise $\epsilon\sim\mathcal{N}(0,\kappa^{2}\mathbf{}{I})$, forming a noisy input $x_{T}=y+\epsilon$. The Student network $f_{\phi}$ predicts a latent representation as $x_{0}=f_{\phi}(x_{T},y,T)$. This output from the Student network is concatenated in the latent space with a Gaussian noise $\epsilon^{\prime}\sim\mathcal{N}(0,\kappa^{2}\mathbf{}{I})$, creating the input for the Downsampler as $\text{concat}(x_{0},\epsilon^{\prime})$. The Downsampler Network $G$ then receives this concatenated latent representation and predicts a reconstructed low-resolution image $\hat{y}$:
$\hat{y}=G(\text{concat}(x_{0},\epsilon^{\prime}))$ (13)
This formulation captures the interplay between the Student Network and the Downsampler Network, where the Student Network transforms the low-resolution input into a high-resolution output, which the Downsampler then processes to recover the original low-resolution distribution. The concatenation mechanism facilitates communication between the two diffusion networks, ultimately enabling consistency in the low-resolution image reconstruction and the computation of the corresponding loss.
2 Ablation Study
2.1 Diffusion Downsampler
The comparison between constructing the downsampler architecture using a diffusion-based approach and using convolutional layers is presented in Table 5. Results marked with O* represent those obtained with the diffusion-based downsampler. As shown in the table, the diffusion-based approach demonstrates a performance improvement. Notably, our study is the first to adopt a diffusion-based downsampler that inherently learns through a downsampling task while achieving upsampling as an outcome. This novel approach underscores the significance of our work in advancing the understanding and application of diffusion-based architectures.
Loss Datasets
RealSR RealSet65
Distill HR LR CLIPIQA(↑) MUSIQ(↑) CLIPIQA(↑) MUSIQ(↑)
O* O* O* 0.7240 63.2628 0.7428 64.0630
[…]
O X O 0.6887 61.6081 0.7076 63.3870
O X X 0.6221 58.7344 0.6564 60.9495
[…]
2.2 Diverse Degradations
Our diffusion-based downsampler enables the generation of diverse low-resolution images from a single high-resolution input, producing a range of downsampling kernels that capture varied degradation patterns. By adjusting the noise input in our latent-space-trained downsampler, we generated multiple low-resolution images to demonstrate this diversity. For visual clarity, these low-resolution images have been upscaled and are displayed in Fig 15. Although these representations are feature-level outputs in the latent space and may differ from actual low-resolution images, they effectively showcase the range of degradation patterns learned by the diffusion-based downsampler. This diversity highlights the downsampler’s capability to model complex and realistic degradation characteristics.